QUESTION 1

Turtle Ltd supplied the following information for the past year:   

N$

Direct material used 50 000

Indirect material used 10 000

Direct labour cost    100 000

Indirect labour cost    25 000

Sales commission paid 17 000

Depreciation: plant and equipment    50 000

Sundry factory overheads 50 000

Depreciation: office building    10 000

Sales 410 000


You are required to determine the following:

1. Prime costs

2. Manufacturing overheads costs

3. Conversion costs

4. Total manufacturing costs

5. Period costs

6. Product costs

7. Non-manufacturing costs
Answer #1

1) Prime cost = 50000+100000 = 150000

2) Manufacturing overhead cost = 10000+25000+50000+50000 = 135000

3) Conversion Cost = 135000+100000 = 235000

4) Total manufacturing cost = 150000+135000 = 285000

5) Period Cost = 17000+10000 = 27000

6) Product cost = 285000

7) Non manufacturing cost = 27000